# Transient Dynamical Phase Diagram of the Spin–Boson Model at Finite Temperature

### v1

@@ I Introduction
One of the most foundational models in the study of quantum open systems is the spin–boson model (SBM) [1]. The SBM comprises a single spin linearly coupled to a bosonic continuum with some spectral density. This model and its extensions have been studied in various contexts going back to the 1960s. Some examples include its connection to the Kondo model, which is similar except in having a fermionic continuum [2]; its applications to chemical dynamics in the presence of a dissipative environment [3, 4], as well as electronic degrees of freedom coupled to optical cavities [5]; and as a paradigm for non-Markovianity in quantum information theory [6].
@@
The SBM has numerous experimental implications and realizations, with some examples of recent interest coming on the one hand from using it to understand dissipation in certain types of qubits [7], and on the other hand from using superconducting circuits [8, 9] or ultracold trapped ions [10] to simulate it.
@@
A particularly intriguing aspect of the SBM is the existence of a quantum phase transition (QPT) between the localized and delocalized phase when the low-energy behavior of the continuum spectral density obeys certain restrictions [1]. The spectral density can be parametrized by a spectral density $J(\omega)$ for $\omega>0$. It is then often useful to consider power-law scaling of the spectral density at low energies, in the sense that $J(\omega)\propto\omega^{s}$ in some range of frequencies $0\leq\omega\lesssim\omega_{\text{c}}$ where $\omega_{\text{c}}$ is a cutoff frequency. At high energies, the spectral density often follows an exponential cut-off, $J(\omega)\propto e^{-\omega/\omega_{c}}$. The case $0<s<1$ is then known as the sub-Ohmic regime, and will be the focus of the present study.
The critical behavior and phase diagram of the zero-temperature sub-Ohmic SBM in equilibrium is known. The infinite bandwidth limit can be understood from analytical considerations [1]. More generally, our understanding is mainly due to pioneering numerical work including a combination of early Monte Carlo studies [11, 12, 13] based on the Anderson–Yuval–Hamann theory [14, 15] and a sequence of studies from the numerical renormalization group [16, 17, 18, 19, 20, 13, 21, 20].
In other works, various numerical methods [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] have been used to study the SBM in terms of properties such as the long-time limit of its relaxation to equilibrium from an initial nonequilibrium state [32]; or its transient response properties after a quench [40, 41, 42, 43]. This view corresponds more closely to the time-dependent measurements performed in some of the experiments [9, 10]. It also enables the study of dynamical crossovers between phases with different coherence characteristics. Several investigations of the QPT have explored its dependence on the temperature of the bath from the dynamical point of view [18, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. In contrast, the temperature dependence of the coherence crossover is not fully understood. The dynamics in general [54] and the crossover in particular [55] have been addressed, but most work has focused on zero temperature [56, 32, 40, 42, 43].
Recently [57], we mapped out the full dynamical transient phase diagram of the QPT and the coherence crossover in the low-temperature limit of the sub-Ohmic SBM using numerically exact short-time dynamics obtained from the inchworm Monte Carlo technique [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90]. However, the experimental importance of considering finite-temperature behavior and the approach to the zero-temperature limit is clear. Therefore, in the present work, we explore the dependence of the dynamical phase diagram on temperature.
@@ II Model
Here we choose the sub-Ohmic functional form where $\alpha$ controls the system–bath interactions, $\omega_{c}$ is the cutoff frequency, and $0<s<1$ is the sub-Ohmic exponent. As in Ref. 57 we use $\omega_{c}=10\Delta$ throughout. For the initial condition, we consider the bath to be decoupled from the spin subsystem and the total density matrix takes the factorized form $\rho_{0}=|1\rangle\langle 1|\otimes e^{-\beta H_{b}}/Z_{b}$ where $\beta=1/T$ and $Z_{b}=\textnormal{Tr}\left\{e^{-\beta H_{b}}\right\}$. Note that we set $\hbar=k_{B}=1$ throughout.
@@
In Ref. 57, we explored the behavior of the model near zero temperature ($T\rightarrow 0$) with a focus on the localization and coherence behavior of the population as a function of coupling strength ($\alpha$) and sub-Ohmic exponent ($s$). In the present work, we turn our attention to extending this analysis to finite temperatures.
@@ III.3 Bath correlation functions
$\times\left[\coth(\frac{\beta\omega}{2})-\frac{2}{\beta\omega}\right]\cos(\omega\tau)d\omega$
@@ III.4 Fitting and Analysis
Following Ref. 57, we fit the dynamics of the subsystem $\langle\sigma_{z}(t)\rangle$ to the functional form
@@ IV.1 Time evolution of the spin polarization
As we noted in Sec. III, we next fit these dynamics with a simple function Eq. (11) in order to extract physically meaningful characteristics, including the long-time offset coefficient ($c$), the renormalized coherence frequency ($\Omega$), and the damping coefficient ($\gamma_{1}$). This will enable us to present a wider view of how the dynamics depends on the various parameters, as well as a simpler picture of the physics. The following subsections will therefore discuss the fitting parameters obtained in this way rather than the time dependence itself. We show plots of additional fit parameters pertaining to the non-oscillating decay component of the evolution ($\gamma_{2}$ and $b$) in Appendix A. For completeness, we include a data file with values of all fit parameters in the Supplemental Material [111].
@@ IV.3.1 Coupling dependence
As expected, coherence generally decreases when the temperature rises. This applies to both decoherence mechanisms: the $\Omega/\Delta=0$ transition and the $\Omega/\gamma_{1}=1$ crossover both occur at lower values of $\alpha$ for all values of $s$. The temperature dependence of the frequency-driven incoherence transition is particularly noteworthy. Although at low temperatures this transition can only be observed in our data for $s\gtrsim 0.5$, at higher temperatures all values of $s$ exhibit a sharp drop to zero frequency within the range of couplings accessible to our methodology. The comparison of the critical $\alpha$ values for the $\Omega/\Delta=0$ transition and the $\Omega/\gamma_{1}=1$ crossover will be discussed in more detail in Sec. IV.4.
We now turn our attention to the coherent side of the transition (i.e. the $\Omega/\Delta>0$ regime). In the left panels of Fig. 4, particularly for small $s$, we observe an initial increase in $\Omega$ at lower values of $\alpha$ followed by a sudden drop to zero at higher values of $\alpha$. This behavior becomes more pronounced as the temperature increases.
@@ V Conclusions
In summary, we explored how temperature reshapes the dynamical phase diagram of the sub‑Ohmic SBM. We found that it weakens localization, amplifies damping, and lowers coherence, maintaining the distinction between the two different decoherence mechanisms we considered. These findings could be potentially useful in guiding near‑term experiments that emulate spin–boson physics in superconducting circuits, trapped ions, or molecular nanomagnets. More broadly, they embody an illuminating proof-of-concept for exploring the dynamical phase diagrams of quantum many-body systems and pose an interesting challenge to numerically exact techniques for simulating open quantum systems away from equilibrium. The question of how—and perhaps in some cases whether—the transient phase diagram eventually decays to the equilibrium physics in the long-time limit remains open.
@@ Acknowledgments
We thank Jianshu Cao, Jan von Delft, Carlos González-Gutiérrez, Olivier Parcollet, Nikolay Prokof’ev, and Andreas Weichselbaum for inspiring discussions. O.G. is supported by the NSF under Grants No. PHY-2441282, PHY-2112738 and OSI-2328774. M.G. is supported by by the Israel Science Foundation (ISF) and the Directorate for Defense Research and Development (DDR&D) grant No. 3427/21, by the ISF grant No. 1113/23 and by the US-Israel Binational Science Foundation (BSF) Grant No. 2020072. G.C. is supported by the ISF (Grant No. 2902/21) and by the PAZY foundation (Grant No. 318/78). We acknowledge high-performance computing support of the R2 compute cluster (DOI: 10.18122/B2S41H) provided by Boise State University’s Research Computing Department, the chimera cluster provided by UMass Boston Research Computing, and the Unity cluster. G.C. and HT. C. are supported by the TAU–ND Joint Research Program funded by the Schlindwein Family.

### v2

@@ I Introduction
One of the most foundational models in the study of open quantum systems is the spin–boson model (SBM) [1]. The SBM comprises a single spin linearly coupled to a bosonic continuum with some spectral density. This model and its extensions have been studied in various contexts going back to the 1960s. Some examples include its connection to the Kondo model, which is similar except in having a fermionic continuum [2]; its applications to chemical dynamics in the presence of a dissipative environment [3, 4], as well as electronic degrees of freedom coupled to optical cavities [5]; and as a paradigm for non-Markovianity in quantum information theory [6].
@@
The SBM has numerous experimental implications and realizations, with some examples of recent interest coming on the one hand from using it to understand dissipation in certain types of qubits [7], and on the other hand from using superconducting circuits [8] or ultracold trapped ions [9] to simulate it.
@@
A particularly intriguing aspect of the SBM is the existence of a quantum phase transition (QPT) between the localized and delocalized phase when the low-energy behavior of the continuum spectral density obeys certain restrictions [1]. The spectral density can be parametrized by a continuous function $J(\omega)$ for $\omega>0$, often with power-law scaling at low energies, in the sense that $J(\omega)\propto\omega^{s}$ in some range of frequencies $0\leq\omega\lesssim\omega_{\text{c}}$ where $\omega_{\text{c}}$ is a cutoff frequency. At high energies, the spectral density often follows an exponential cut-off, $J(\omega)\propto e^{-\omega/\omega_{c}}$. The Ohmic SBM ($s=1$) features a Berezinskii–Kosterlitz–Thouless transition between the localized and delocalized states, and the corresponding critical coupling, as well as the critical coupling for the onset of coherence (Toulouse point) are analytically known [1, 4]. The case $0<s<1$ is known as the sub-Ohmic regime, and will be the focus of the present study. This regime is of particular interest, on the one hand due to its rich phase diagram with a second order localization QPT and regimes of coherence and decoherence, and on the other hand due to its experimental relevance for superconducting and mesoscopic circuits [10, 11, 8, 12, 13] and trapped ion systems [14, 15].
The critical behavior and phase diagram of the zero-temperature sub-Ohmic SBM in equilibrium are known. The infinite bandwidth limit can be understood from analytical considerations [1]. More generally, our understanding rests on the quantum–classical correspondence between the SBM and an Ising spin chain with $1/r^{1+s}$ pair interaction [16, 17, 18, 19]. Within this framework, the localization transition of the SBM is in the same universality class as the classical Ising transition and, therefore, the mapping predicts the structure of its equilibrium phase diagram and its critical scaling [20, 21]. The quantitative characterization of the equilibrium QPT has subsequently been refined by a variety of approaches. Early numerical renormalization group and renormalization group studies initiated a debate regarding the range of validity of the quantum–classical mapping and the associated critical exponents [22, 23, 24]. Later work clarified the origin of the discrepancies [25, 26] and established consistency with the Ising-model predictions by employing improved renormalization-group treatments together with complementary equilibrium methods, including quantum Monte Carlo [27], exact diagonalization [28], density matrix renormalization group approaches [29], variational matrix-product-state calculations [30], and variational simulations with polaron ansatz [31, 32]. As a result, both equilibrium critical couplings and critical exponents have been determined reliably across the full sub-Ohmic regime.
The real-time dynamics in the sub-Ohmic regime has been studied with a variety of approximate and numerically exact methods. Early approximate approaches, including variational ansätze [33], perturbative methods [34], and flow equation calculations [35], established a qualitative picture of coherent oscillations at weak dissipation and incoherent relaxation at stronger dissipation. These treatments also highlighted the renormalization of the tunneling element, its nontrivial temperature dependence, and the pronounced role of low-frequency bath modes in controlling decoherence. More detailed coherence maps and dynamical phase diagrams at zero temperature have been obtained using time-dependent variational wave-function approaches based on Davydov $\mathrm{D}_{1}$ trial states and their multi-$\mathrm{D}_{1}$ extensions [36, 37, 38, 39].
Beyond these approximate approaches, a number of numerically exact techniques have been applied to the dynamics of the SBM in general, including its sub-Ohmic regime [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. They provide complementary perspectives and benchmarks on transient relaxation, long-time equilibration, and coherence properties of the model. The dynamical view corresponds more closely to the time-dependent measurements performed in some of the experiments [8, 9]. Interestingly, these studies indicate that the equilibrium localization transition and the loss of coherent oscillations in the dynamics need not coincide. Coherent oscillations can persist even in regions of the parameter space that are localized in equilibrium, while incoherent decay at zero temperature occurs only above a critical sub-Ohmic exponent $s_{c}$. There is some disagreement regarding the precise value of $s_{c}$, but it is generally found to be in the range 0.40.50.4-0.5 [63, 49, 64, 38, 52, 61, 39, 62].
These transitions can be illustrated by sketching a transient dynamical phase diagram [65, 62]. Early numerical work by Nalbach and Thorwart [65] employed the quasiadiabatic propagator path integral (QUAPI) approach and showed that the low‑frequency bath modes generate a dynamical asymmetry, leading to a transient quasiequilibrium and ultraslow relaxation. Note that these dynamical features of the sub-Ohmic SBM are unique to the initial condition where the state is factorized, but the bath is allowed to relax to an equilibrium state congruent with the initial state of the spin, which we will refer to as the shifted bath case. Another common choice is the unshifted bath initial condition, where the state is factorized but the bath is at thermal equilibrium by itself. The dynamics of the sub-Ohmic SBM has been studied for both the shifted [24, 34, 66, 65] and unshifted [44, 45, 52, 61, 67, 62] initial conditions, including works that discuss both choices [46, 36, 37, 38, 49, 64, 39]. Coherence properties of the model in general are highly sensitive to the choice of initial preparation.
Recently [62], we mapped out the complete dynamical transient phase diagram of the QPT and the coherence regions of the sub-Ohmic SBM with the unshifted bath initial condition. We used numerically exact short-time dynamics obtained from the inchworm Monte Carlo technique [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 75, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100] and focused on the low-temperature limit. The temperature dependence of the SBM dynamics in general has been studied in different contexts [101, 65, 102, 64, 103, 104, 105, 67, 106, 107, 108, 66]. For the shifted initial condition [65], increasing temperature suppresses the dynamical asymmetry, and the incoherence crossover temperature increases with decreasing $s$. However, for the unshifted initial condition, the temperature dependence of the transient dynamical phase diagram has not yet been explored. Here we extend the dynamical phase diagram to the finite temperature regime.
@@ II Model
Here we choose the sub-Ohmic functional form where $\alpha$ controls the system–bath interactions, $\omega_{c}$ is the cutoff frequency, and $0<s<1$ is the sub-Ohmic exponent. As in Ref. 62 we use $\omega_{c}=10\Delta$ throughout. For the initial condition, we consider the bath to be decoupled from the spin subsystem and the total density matrix takes the factorized form $\rho_{0}=|1\rangle\langle 1|\otimes e^{-\beta H_{b}}/Z_{b}$ where $\beta=1/T$ and the bath configuration is chosen to be unshifted $Z_{b}=\textnormal{Tr}\left\{e^{-\beta H_{b}}\right\}$. Note that we set $\hbar=k_{B}=1$ throughout.
@@
In Ref. 62, we explored the behavior of the model near zero temperature ($T\rightarrow 0$) with a focus on the localization and coherence behavior of the population as a function of coupling strength ($\alpha$) and sub-Ohmic exponent ($s$). In the present work, we turn our attention to extending this analysis to finite temperatures.
@@ III.3 Bath correlation functions
$\times\left[\coth(\frac{\beta\omega}{2})-\frac{2}{\beta\omega}\right]\cos(\omega\tau)d\omega.$
@@ III.4 Fitting and Analysis
Following Ref. 62, we fit the dynamics of the subsystem $\langle\sigma_{z}(t)\rangle$ to the functional form
@@ IV.1 Time evolution of the spin polarization
As we noted in Sec. III, we next fit these dynamics with a simple function Eq. (11) in order to extract physically meaningful characteristics, including the long-time offset coefficient ($c$), the renormalized coherence frequency ($\Omega$), and the damping coefficient ($\gamma_{1}$). This will enable us to present a wider view of how the dynamics depends on the various parameters, as well as a simpler picture of the physics. The following subsections will therefore discuss the fitting parameters obtained in this way rather than the time dependence itself. We show plots of additional fit parameters pertaining to the non-oscillating decay component of the evolution ($\gamma_{2}$ and $b$) in Appendix A. For completeness, a data file with values of all fit parameters is openly available [128].
@@ IV.3.1 Coupling dependence
As expected, coherence generally decreases when the temperature rises. This applies to both decoherence mechanisms: the $\Omega/\Delta=0$ transition and the $\Omega/\gamma_{1}=1$ crossover both occur at lower values of $\alpha$ for all values of $s$.
The temperature dependence of the frequency-driven incoherence transition is particularly noteworthy. In the left panels of Fig. 4, we observe for some values of $s$ a sudden drop of the frequency to zero at higher values of $\alpha$. Although at low temperatures the transition to $\Omega=0$ can only be observed in our data for $s\gtrsim 0.5$, at higher temperatures all values of $s$ exhibit a sharp drop to $\Omega=0$ within the range of couplings accessible to our methodology. This observation may explain the uncertainty in the literature about the critical sub-Ohmic exponent $s_{c}$ at low temperature [63, 49, 64, 38, 52, 61, 39]. The comparison of the critical $\alpha$ values for the $\Omega/\Delta=0$ transition and the $\Omega/\gamma_{1}=1$ crossover will be discussed in more detail in Sec. IV.4.
We also observe, in particular for smaller values of $s$, that in some cases an initial increase in $\Omega$ can be seen at lower values of $\alpha$. At small values of $s$ and higher temperatures, this precedes the drop to $\Omega=0$. The existence of the transition for $s\lesssim 0.5$ appears in conjunction with this nonmonotonic dependence of $\Omega$ with respect to $\alpha$.
@@ V Conclusions
In summary, we explored how temperature reshapes the dynamical phase diagram of the sub‑Ohmic SBM with unshifted bath initial conditions. We found that it weakens localization, amplifies damping, and lowers coherence, maintaining the distinction between the two different decoherence mechanisms we considered. These findings could be potentially useful in guiding near‑term experiments that emulate spin–boson physics in superconducting circuits, trapped ions, or molecular nanomagnets. More broadly, they embody an illuminating proof-of-concept for exploring the dynamical phase diagrams of quantum many-body systems and pose an interesting challenge to numerically exact techniques for simulating open quantum systems away from equilibrium. The question of how—and perhaps in some cases whether—the transient phase diagram eventually decays to the equilibrium physics in the long-time limit remains open.
@@ Acknowledgments
We thank Jianshu Cao, Jan von Delft, Carlos González-Gutiérrez, Olivier Parcollet, Nikolay Prokof’ev, and Andreas Weichselbaum for inspiring discussions. O.G. is supported by the NSF under Grants No. PHY-2441282, PHY-2112738 and OSI-2328774. M.G. is supported by by the Israel Science Foundation and the Directorate for Defense Research and Development (DDR&D) grant No. 3427/21, by the Israel Science Foundation grant No. 1113/23 and by the US-Israel Binational Science Foundation (BSF) Grant No. 2020072. G.C. is supported by the Israel Science Foundation (Grant No. 2902/21) and by the PAZY foundation (Grant No. 318/78). We acknowledge high-performance computing support of the R2 compute cluster (DOI: 10.18122/B2S41H) provided by Boise State University’s Research Computing Department, the chimera cluster provided by UMass Boston Research Computing, and the Unity cluster. G.C. and HT. C. are supported by the TAU–ND Joint Research Program funded by the Schlindwein Family.